# Beyond H&E: Unlocking Pathological Insights with Polarization via Self-supervised Learning

### v1

@@ Abstract
Histopathology image analysis is fundamental to digital pathology, with hematoxylin and eosin (H&E) staining as the gold standard for diagnostic and prognostic assessments. While H&E imaging effectively highlights cellular and tissue structures, it lacks sensitivity to birefringence and tissue anisotropy, which are crucial for assessing collagen organization, fiber alignment, and microstructural alterations—key indicators of tumor progression, fibrosis, and other pathological conditions. To bridge this gap, we propose PolarHE, a dual-modality fusion framework that integrates H&E with polarization imaging, leveraging the latter’s ability to enhance tissue characterization. Our approach employs a feature decomposition strategy to disentangle common and modality-specific features, ensuring effective multimodal representation learning. Through comprehensive validation, our approach significantly outperforms previous methods, achieving an accuracy of 86.70% on the Chaoyang dataset and 89.06% on the MHIST dataset. Moreover, polarization property visualization reveals distinct optical signatures of pathological tissues, highlighting its diagnostic potential. t-SNE visualizations further confirm our model effectively captures both shared and unique modality features, reinforcing the complementary nature of polarization imaging. These results demonstrate that polarization imaging is a powerful and underutilized modality in computational pathology, enriching feature representation and improving diagnostic accuracy. PolarHE establishes a promising direction for multimodal learning, paving the way for more interpretable and generalizable pathology models. Our code will be released after paper acceptance.
Keywords:
Histopathology Polarization Imaging Multimodal Fusion.
1 Introduction
@@
Histopathology image analysis is fundamental in digital pathology, with hematoxylin and eosin (H&E) staining serving as the gold standard for diagnosing and prognosticating various diseases, particularly cancer. While H&E imaging effectively highlights cellular and tissue structures, it lacks the ability to capture birefringence and tissue anisotropy, which provide critical insights into collagen organization, fiber alignment, and microstructural alterations—key indicators of tumor progression, fibrosis, and other pathological conditions [6, 26].
Polarization imaging, which leverages the fundamental properties of light, has emerged as a promising auxiliary modality for histopathological analysis. By capturing tissue birefringence, scattering properties, and depolarization effects, it provides additional contrast that reveals microstructural and biochemical changes beyond what is visible in standard H&E staining [8]. Studies have demonstrated its effectiveness in cervical intraepithelial neoplasia grading and breast carcinoma classification, enabling high-accuracy distinction of pathological features such as collagen alignment and nuclear organization [5, 6, 10]. These advantages position polarization imaging as a powerful tool for early disease detection and precise tissue characterization.
@@
Despite its potential, polarization imaging remains underutilized in computational pathology, primarily due to the lack of established methodologies for its integration with conventional H&E-stained slides. Recent advancements in self-supervised learning (SSL) present a promising avenue for effectively incorporating multimodal data [17]. SSL frameworks, which learn meaningful representations from unlabeled data by solving pretext tasks, have shown remarkable success across various domains, including medical imaging [11, 13, 23]. By jointly learning from H&E and polarization imaging in a self-supervised manner, it becomes possible to develop robust representations that capture both shared and modality-specific features, enhancing downstream tasks such as patch classification and tissue segmentation [21].
@@
Therefore, we propose PolarHE, a dual-modality fusion framework that integrates H&E and polarization imaging for enhanced representation learning in pathology. Our framework employs a feature decomposition strategy to disentangle common and modality-specific features, ensuring shared features remain aligned across modalities while preserving the discriminative properties of unique modality features. We validate our framework on a patch-level classification task, where PolarHE significantly outperforms previous methods. Furthermore, polarization property visualization confirms polarization captures distinct optical tissue signatures, reinforcing the complementary role of polarization imaging in pathology. t-SNE visualizations further highlight the effectiveness of our model in learning meaningful multimodal representations. Our key contributions are as follows: 1) We introduce PolarHE, a novel dual-modality fusion framework combining H&E and polarization imaging for enhanced histopathology analysis; 2) We adopt a feature decomposition strategy that disentangles common and modality-specific features, facilitating multimodal learning; 3) We demonstrate model effectiveness through extensive experiments and visualizations, showing significant performance improvements on patch classification tasks.
2 Feature Decomposition for Multimodal Fusion
@@
2.1 Cross-correlation Matrix for Feature Dependency Analysis
To measure dependencies between embedding dimensions, we compute cross-correlation matrices. Given two embedding vectors $F^{A},F^{B}\in\mathbb{R}^{K}$, their correlation matrix $\mathcal{C}$ is defined as:
@@ 2.1 Cross-correlation Matrix for Feature Dependency Analysis
2.2 Cross-modal Representation Decoupling
To ensure proper feature disentanglement, we partition the embedding space $K$ into: $K_{c}$ (common representations) and $K_{u}$ (unique representations) with $K_{c}+K_{u}=K$. In Figure 1, the common representations are identical (highlighted in red) and encode modality-invariant features, while the modality-unique representations are decorrelated (highlighted in purple) and capture specialized modality-dependent structures.
@@ 2.2 Cross-modal Representation Decoupling
2.3 Intra-modal Representation Enhancement
To reinforce feature stability and prevent collapse of decoupled unique dimensions, we introduce intra-modal training that covers all the embedding dimensions, ensuring that the model learns robust within-modality representations. For each modality, a cross-correlation matrix $\mathcal{C_{\text{H}}}$ (or $\mathcal{C_{\text{P}}}$) is generated from the full dimensions of the embedding vectors ${F^{H}_{1}}$ and ${F^{H}_{2}}$ (or ${F^{P}_{1}}$ and ${F^{P}_{2}}$), optimizing:
@@ 2.3 Intra-modal Representation Enhancement
3 Experiments
3.1 Polarization System and Paired Dataset Construction
Polarization Imaging System for Pathology. To facilitate multimodal pathology image analysis, we construct a polarization imaging system for data acquisition and processing, as illustrated in Figure 2. This system is designed to capture full-vector polarization imaging, enabling comprehensive tissue characterization beyond conventional H&E staining. We employ a 635nm red light source as the incident light and perform full-field scanning at 4× magnification to ensure uniform tissue coverage across all pathology slides. Each field of view undergoes polarization measurement, generating a 16-dimensional polarization matrix $M_{s(u,v)}$ for each pixel $(u,v)$. This results in a 16-channel polarization representation, providing rich structural and birefringence information, which is essential for analyzing tissue anisotropy and microstructural organization.
Figure 2: Polarization Imaging Acquisition and Processing System.
PolarHE Paired Dataset Collection. To establish a precisely aligned multimodal dataset, we store the corresponding H&E-stained whole-slide images (WSIs) alongside the polarization images, preserving their structural and morphological details. To enable direct pixel-wise comparison, we perform image registration between the two modalities, ensuring precise spatial alignment. This results in a dataset of 26 paired PolarHE WSIs, from which we extract 4035 multimodal image patches of size 2304 × 1296. These patches are further cropped into 224 × 224 regions for model input, providing a rich and diverse dataset that enables SSL and downstream pathology analysis. This dataset facilitates the joint learning of common and unique representations across imaging modalities, making it an ideal benchmark for computational pathology.
@@ 3.1 Polarization System and Paired Dataset Construction
Implementation. Each branch of PolarHE consists of a dedicated backbone and a three-layer MLP projector, each with an output dimension of 8192. We adopt the UNI ViT-L/16 encoder [3] as the backbone. Training is performed on embedding representations obtained after the projector, where the feature space is partitioned into common and unique components, with 75% allocated to common features. The backbone is subsequently transferred to downstream tasks, while the trade-off parameter ($\lambda$) for the loss terms is set to 0.0051 [28].
3.2 Downstream Patch Classification
@@
Datasets. To assess the effectiveness of PolarHE, we conduct patch classification experiments on two benchmark histopathology datasets: Chaoyang Dataset [29] consists of colon tissue slides collected from Chaoyang Hospital. Each patch is 512 × 512 pixels, with the dataset split as follows: Training Set: 1,111 normal, 842 serrated, 1,404 adenocarcinoma, and 664 adenoma patches; Testing Set: 705 normal, 321 serrated, 840 adenocarcinoma, and 273 adenoma patches. MHIST Dataset [22] consists of 2,175 training images and 977 testing images from 328 WSIs. These images are labeled into two classes: Hyperplastic Polyps (HPs) and Sessile Serrated Adenomas (SSAs). From these WSIs, 3,152 diagnostically relevant patches of size 224 × 224 pixels are extracted,representing clinically significant tissue regions.
Patch Classification by Linear Probing. To validate the effectiveness of our pretrained model, we perform linear probing [11] on the Chaoyang and MHIST datasets, demonstrating that PolarHE’s learned representations are highly discriminative, requiring minimal task-specific fine-tuning. Our method, PolarHE, achieves state-of-the-art performance on both datasets. On the Chaoyang Dataset, it achieves an accuracy of 86.70%, an F1 score of 86.42%, and an AUC of 95.96%. On the MHIST Dataset, it achieves an accuracy of 89.06%, an F1 score of 88.97%, and an AUC of 94.10%. These results, presented in Table 1, demonstrate that PolarHE significantly outperforms existing methods, validating the effectiveness of our dual-modality fusion.
Table 1: Test Performance on Chaoyang [29] and MHIST [22] datasets.
Chaoyang Dataset MHIST Dataset
Method Acc F1 AUC Method Acc F1 AUC
DINO [1] 80.2 - - DINO [1] 83.10 81.60 89.78
HSA-NRL [29] 83.40 76.54 94.51 BarlowTwins [28] 81.27 - -
SSAT [4] 82.52 - - MoCo-v2 [11] 85.88 - -
ViT-DAE [24] 80.12 78.60 - Bitfit [27] 82.40 80.86 89.60
INCV [2] 80.34 74.11 92.63 SSF [15] 82.29 80.85 89.29
2HSL [7] 82.90 - - RMSGD [9] 82.58 - -
OUSM [25] 80.53 73.70 93.69 TPP [14] 83.42 82.21 91.38
SELF [18] 80.49 75.31 93.99 DiRL [12] 78.20 - 87.1
DiRL [12] 83.60 - 95.0 Virchow [19] 83.40 83.5 -
SH-PEFT [16] 84.80 80.6 - CTransPath [20] 83.75 - -
UNI [3] 85.51 85.29 95.21 UNI [3] 85.63 85.57 91.98
PolarHE (ours) 86.70 86.42 95.96 PolarHE (ours) 89.06 88.97 94.10
@@
Ablation Study. Table 2 shows removing intra-modal training or modality decoupling reduces accuracy, confirming their importance. Without intra-modal training, performance drops, especially on Chaoyang, highlighting its role in reinforcing modality-specific features. Without modality decoupling, accuracy also declines, showing the need for feature separation. The full model performs best, validating both components. For common representation ratio, the downstream performance increases and decreases smoothly along with the change of the percentage of common dimensions, which indicates the sparsity of the common embedding space.
Table 2: Ablation study on loss components and common representation ratio. Intra and decoup. correspond to intra-modal training and modality decoupling. We report the Accuracy Index in both experiments.
@@
3.3 Polarization Property Visualization
Polarization imaging introduces a contrast mechanism for visualizing tissue microstructure and biochemical properties, enabling precise differentiation of pathological features. Unlike H&E staining, where collagen fibers appears pink (indistinguishable from muscle and blood vessels), the fast-axis orientation effectively highlights collagen organization in full-vector polarization imaging, aiding in structural differentiation.
Similarly, endometrioid carcinoma exhibits unique polarization signatures. Compared to normal hepatocytes, cancerous regions show increased DNA replication activity, higher nuclear stiffness, and stronger depolarization signals. The difference in nuclear depolarization saturation provides a potential marker for distinguishing cancerous regions, assessing tumor differentiation, and evaluating malignancy progression. These findings suggest that polarization imaging provides valuable tissue biomarkers, offering enhanced diagnostic capabilities beyond conventional H&E analysis.
3.4 T-SNE Visualization of Common and Unique Representations
To further investigate the representations learned by PolarHE, we employ t-SNE visualization to analyze the distribution of common and unique features: Shared features between polarization and H&E images exhibit strong alignment in feature space, confirming effective cross-modal feature learning; Unique modality-specific features remain well-separated, demonstrating that our model successfully preserves discriminative modality-dependent information. The distinct separation of common and unique features highlights the effectiveness of PolarHE in learning robust multimodal features, reinforcing the complementary role of polarization imaging in pathology analysis.
Figure 4: t-SNE Visualization (Red: H&E Common Feature; Blue: Polarization Common Feature; Green: H&E Unique Feature; Purple: Polarization Unique Feature.)
4 Conclusion
We propose PolarHE, a dual-modality fusion framework that integrates polarization imaging with H&E staining for enhanced representation learning in pathology. 1By leveraging a feature decomposition strategy, our approach effectively disentangles common and modality-specific features, leading to improved multimodal representation learning. Through comprehensive experiments, we demonstrated that the integration of H&E and polarization representations leads to significantly better performance compared to existing methods. Additionally, t-SNE visualizations confirm that our model successfully aligns shared features while preserving modality-specific distinctions, reinforcing the complementary nature of polarization imaging. These findings underscore the untapped potential of polarization imaging as a powerful auxiliary modality in computational pathology. By enhancing diagnostic accuracy and model robustness, PolarHE marks a promising step toward more interpretable and automated pathology analysis, paving the way for advancements in multimodal deep learning for clinical applications.

### v2

# Beyond H&E: Unlocking Pathological Insights with Polarization Imaging

@@ Abstract
Histopathology image analysis is fundamental to digital pathology, with hematoxylin and eosin (H&E) staining as the gold standard for diagnostic and prognostic assessments. While H&E imaging effectively highlights cellular and tissue structures, it lacks sensitivity to birefringence and tissue anisotropy, which are crucial for assessing collagen organization, fiber alignment, and microstructural alterations—key indicators of tumor progression, fibrosis, and other pathological conditions. To bridge this gap, we construct a polarization imaging system and curate a new dataset of over 13,000 paired Polar-H&E images. Visualizations of polarization properties reveal distinctive optical signatures in pathological tissues, underscoring its diagnostic value. Building on this dataset, we propose PolarHE, a dual-modality fusion framework that integrates H&E with polarization imaging, leveraging the latter’s ability to enhance tissue characterization. Our approach employs a feature decomposition strategy to disentangle common and modality-specific features, ensuring effective multimodal representation learning. Through comprehensive validation, our approach significantly outperforms previous methods, achieving an accuracy of 86.70% on the Chaoyang dataset and 89.06% on the MHIST dataset. These results demonstrate that polarization imaging is a powerful and underutilized modality in computational pathology, enriching feature representation and improving diagnostic accuracy. PolarHE establishes a promising direction for multimodal learning, paving the way for more interpretable and generalizable pathology models.
I Introduction
@@
Histopathology image analysis is fundamental in digital pathology, with hematoxylin and eosin (H&E) staining serving as the gold standard for diagnosing and prognosticating various diseases, particularly cancer [xiang2025moc]. While H&E imaging highlights cellular and tissue structures, it lacks the ability to capture birefringence and tissue anisotropy, which provide critical insights into collagen organization, fiber alignment, and microstructural alterations—key indicators of tumor progression, fibrosis, and other pathological conditions [dong2021polarization, zhang2024incidence, zheng2025diffusion].
Polarization imaging, which leverages the fundamental properties of light, has emerged as a promising auxiliary modality for histopathological analysis. By capturing tissue birefringence, scattering properties, and depolarization effects, it provides additional contrast that reveals microstructural and biochemical changes beyond what is visible in standard H&E staining [he2021polarisation]. More specifically, polarization imaging enables non-invasive acquisition of birefringence characteristics and sub-resolution structural information by analyzing how biological tissues modulate the polarization state of incident light waves—typically quantified through Stokes vectors, Mueller matrices, and derived optical parameters [hao2024cartesian, he2021polarisation]. This imaging mechanism demonstrates high sensitivity to features such as collagen fiber alignment and cell membrane integrity, which are critical indicators of pathological processes.
Recent studies have validated the clinical potential of polarization imaging across diverse applications, including tumor boundary delineation, fibrosis evaluation, cervical intraepithelial neoplasia grading, and breast carcinoma classification [wang2024pathology, pati2024Accelerating, dong2021Deriving, dong2021polarization, hou2022polarimetry]. By enabling high-accuracy distinction of pathological features such as collagen alignment and nuclear organization, polarization imaging emerges as a powerful complement to traditional histopathology—particularly for early disease detection and precise tissue characterization.
Despite its potential, polarization imaging remains underutilized in computational pathology, primarily due to the lack of established datasets and methodologies for its integration with conventional H&E-stained slides. To bridge this gap, we construct a polarization imaging system and collect a novel dataset of paired H&E-stained and polarization images. It comprises over 13,000 spatially aligned image patches, capturing both conventional histological structure and polarization-based birefringence features. This dataset provides a rich foundation for exploring complementary modality representations in tissue analysis. Polarization property visualization confirms polarization captures distinct optical tissue signatures, reinforcing the complementary role of polarization imaging in pathology.
@@
As discussed above, H&E and polarization images offer complementary information for clinical pathology. Effectively leveraging this complementarity is crucial, especially considering that polarization imaging is costly and less accessible in routine clinical practice. To address this, we propose a dual-modality fusion strategy that integrates both H&E and polarization data during training. This enables the model to learn richer, more informative representations, allowing the resulting encoders to generalize better on downstream tasks. Recent advancements in self-supervised learning (SSL) offer a promising avenue for multimodal fusion without requiring extensive manual annotations [nguyen2024towards]. SSL frameworks, which learn meaningful representations from unlabeled data by solving pretext tasks, have shown remarkable success across various domains, including medical imaging [kang2023benchmarking, koohbanani2021self, wu2024voco, dai2023semi]. By jointly learning from H&E and polarization imaging in a self-supervised manner, it becomes possible to develop robust representations that capture both shared and modality-specific features, enhancing downstream tasks such as patch classification and tissue segmentation [wang2024decoupling].
@@
Therefore, we propose PolarHE, a dual-modality fusion framework that integrates H&E and polarization imaging for enhanced representation learning in pathology. Our framework employs a feature decomposition strategy to disentangle common and modality-specific features, ensuring shared features remain aligned across modalities while preserving the discriminative properties of unique modality features. We validate our framework on a patch-level classification task, where PolarHE significantly outperforms previous methods.
Our key contributions are as follows: 1) We construct the first dual-modality Polarization Pathology Dataset, which includes over 13,000 image patches with precise spatial registration between H&E and polarization imaging; 2) We introduce PolarHE, a novel dual-modality fusion framework combining H&E and polarization imaging for enhanced histopathology analysis. We adopt a feature decomposition strategy that disentangles common and modality-specific features, facilitating multimodal learning; 3) We demonstrate model effectiveness through extensive experiments, showing significant performance improvements on patch classification tasks.
II Polarization System and Paired Dataset Construction
Polarization Imaging System for Pathology. To facilitate multimodal pathology image analysis, we construct a polarization imaging system for data acquisition and processing, as illustrated in Figure 1. This system is designed to capture full-vector polarization imaging, enabling comprehensive tissue characterization beyond H&E staining.
The acquisition process begins with the use of specialized slide scanners to obtain Whole-Slide Images (WSIs) of H&E staining. Subsequently, the samples are placed on a stage for polarimetric imaging. We employ a 635nm red light source as the incident light and perform full-field scanning at 4× magnification to ensure uniform tissue coverage across all pathology slides. After autofocus determines the optimal focal plane, the system scans the sample region-by-region with a preset step size. Each field of view undergoes polarization measurement, generating a 16-dimensional polarization matrix $M_{s(u,v)}$ for each pixel $(u,v)$. This results in a 16-channel polarization representation, providing rich structural and birefringence information, which is essential for analyzing tissue anisotropy and microstructural organization.
Given that samples are large and a single shot cannot cover the entire slice, an ITKMontage [zukic2021itkmontage] module is employed for multi-modality image registration and stitching. To address issues such as size consistency and missing data in multi-modality image stitching, the system automatically generates configuration files to record tile position information. It also utilizes boundary detection and virtual tile generation techniques. Additionally, grayscale uniformity correction and brightness normalization are applied. Ultimately, this process results in the formation of complete panoramic WSI images under a 4× objective lens.
Figure 1: Polarization Imaging Acquisition and Processing System.
Spatial Alignment via Image Registration. To establish a precisely aligned multimodal dataset, we store the corresponding H&E-stained WSIs alongside the polarization images, preserving their structural and morphological details. To enable pixel-wise comparison across modalities, we design a multi-stage registration pipeline that ensures precise spatial alignment. Initial pre-processing standardizes the image scales and removes extraneous blank regions to mitigate feature interference. The polarization-derived grayscale image is used as the registration reference due to its structural consistency with other modalities. Feature matching is performed using the SuperPoint algorithm [detone2018superpoint], a deep learning-based detector that extracts cross-modal keypoints for robust correspondence estimation. Rigid registration aligns global structures via affine transformations. Subsequent non-rigid registration refines local deformations using B-spline modeling [unser2002b] combined with curvature regularization [myronenko2006non] to prevent artifacts. After registration, all modalities are resampled to 2304×1296 pixels and aligned with the polarization field-of-view (FOV). Adaptive thresholding is then applied to exclude tissue-free or artifact-dominated regions. The validated FOVs are partitioned into 224×224 pixel patches using a sliding window approach, preserving multi-modality alignment through strict spatial correspondence.
This process yields a multimodal dataset comprising over 13,000 high-quality, spatially aligned patches. This dataset enables integrated learning from H&E and polarization modalities, providing a foundation for investigating multimodal fusion strategies and advancing downstream pathology analysis.
Polarization Property Visualization. Polarization imaging offers a powerful contrast mechanism for visualizing tissue microstructure and biochemical composition, enabling fine-grained differentiation of pathological features. This imaging modality derives pixel-wise optical parameters through Mueller matrix decomposition, capturing subtle light–tissue interactions. The key extracted parameters include:
1) Phase Retardation: quantifies the phase shift between orthogonal polarization components introduced by birefringent structures such as collagen fibers. It serves as an indicator of microstructural anisotropy and fiber density; 2) Fast-axis Orientation: represents the principal axis of birefringence at each pixel, revealing the alignment and directional organization of fibrous tissue components; 3) Depolarization Power: measures the extent to which polarized light becomes randomized after passing through tissue. Higher depolarization is typically associated with structural heterogeneity, such as densely packed or disorganized cell nuclei [yang2024data].
As shown in the left part of Figure 2, compared to traditional H&E staining—where collagen fibers often appear pink and are visually indistinguishable from muscle or blood vessels—polarization imaging, particularly through fast-axis orientation maps, enables clear delineation of collagen structures. This allows for more accurate assessment of fiber arrangement and tissue integrity, which is especially valuable in detecting fibrosis or tumor infiltration.
In the right part of Figure 2, pathological cases such as endometrioid carcinoma exhibit distinctive polarization signatures. Compared to normal hepatocytes, cancerous tissues show elevated DNA replication, increased nuclear stiffness, and stronger depolarization signals, likely reflecting chromatin condensation and cellular heterogeneity. Additionally, cilia located along cavity edges may display polarization-sensitive characteristics. These features indicate variations in depolarization saturation can serve as optical biomarkers for identifying malignant regions and evaluating tumor progression.
Together, these observations underscore the unique capability of polarization imaging to extract interpretable and quantitative optical markers—providing a powerful complement to conventional H&E analysis and substantially enhancing diagnostic resolution in digital pathology.
III Methodology
Overview. Unlike existing multimodal SSL methods that typically operate on loosely or semantically aligned modalities [wang2024decoupling, zbontar2021barlow], our framework is tailored for the histopathology domain, where precise spatial alignment between H&E and polarization images enables meaningful feature-level fusion. While the input modalities are pixel-aligned, we deliberately perform fusion at the representation level to ensure flexibility and generalizability. By exploiting their complementary characteristics—H&E capturing morphology and polarization revealing sub-resolution optical cues—we design a polarization-guided training strategy that enriches H&E representation learning. This allows the H&E encoder to absorb polarization-driven structural cues, improving performance even when only H&E is available at inference. To the best of our knowledge, this is the first SSL framework to jointly learn from H&E and polarization imaging in digital pathology.
@@ III Methodology
Cross-correlation Matrix for Feature Dependency Analysis. To measure dependencies between embedding dimensions, we compute cross-correlation matrices. Given two embedding vectors $F^{A},F^{B}\in\mathbb{R}^{K}$, correlation matrix $\mathcal{C}$ is defined as:
@@
Cross-modal Representation Decoupling. To ensure proper feature disentanglement, we partition the embedding space $K$ into: $K_{c}$ (common representations) and $K_{u}$ (unique representations) with $K_{c}+K_{u}=K$. In Figure 3, the common representations are identical (highlighted in red) and encode modality-invariant features, while the modality-unique representations are decorrelated (highlighted in purple) and capture specialized modality-dependent structures.
@@
Intra-modal Representation Enhancement. To reinforce feature stability and prevent collapse of decoupled unique dimensions, we introduce intra-modal training that covers all the embedding dimensions, ensuring that the model learns robust within-modality representations. For each modality, a cross-correlation matrix $\mathcal{C_{\text{H}}}$ (or $\mathcal{C_{\text{P}}}$) is generated from the full dimensions of the embedding vectors $F^{H}_{1}$} and $F^{H}_{2}$} (or $F^{P}_{1}$} and $F^{P}_{2}$}), optimizing:
@@
IV Experiments
@@
Implementation. Each branch of PolarHE consists of a dedicated backbone and a three-layer MLP projector, each with an output dimension of 8192. We adopt the UNI ViT-L/16 encoder [chen2024towards] as the backbone, as it provides a strong H&E-specific representation baseline, which not only facilitates effective integration with polarization features but also ensures that observed gains can be attributed to our proposed fusion strategy rather than backbone selection. Training is performed on embedding representations obtained after the projector, where the feature space is partitioned into common and unique components, with 75% allocated to common features. During training, both encoder and projector are jointly optimized with a small learning rate of 1e-5. We apply data augmentations including random horizontal/vertical flipping and rotation. The encoder is subsequently transferred to downstream tasks, while the trade-off parameter ($\lambda$) for the loss terms is set to 0.0051.
Downstream Patch Classification. To assess the effectiveness of PolarHE, we conduct patch classification experiments on two benchmark histopathology datasets:
@@
1. Chaoyang Dataset [zhu2021hard] consists of colon tissue slides collected from Chaoyang Hospital. Each patch is 512 × 512 pixels, with the dataset split as follows: Training Set: 1,111 normal, 842 serrated, 1,404 adenocarcinoma, and 664 adenoma patches; Testing Set: 705 normal, 321 serrated, 840 adenocarcinoma, and 273 adenoma patches.
2. MHIST Dataset [wei2021petri] consists of 2,175 training images and 977 testing images from 328 WSIs. These images are labeled into two classes: Hyperplastic Polyps (HPs) and Sessile Serrated Adenomas (SSAs). From these WSIs, 3,152 diagnostically relevant patches of size 224 × 224 pixels are extracted,representing clinically significant tissue regions.
TABLE I: Test performance on Chaoyang [zhu2021hard] dataset.
Method Acc F1 AUC
DINO [caron2021emerging] 80.21 - -
HSA-NRL [zhu2021hard] 83.40 76.54 94.51
SSAT [das2024limited] 82.52 - -
ViT-DAE [xu2023vit] 80.12 78.60 -
INCV [chen2019understanding] 80.34 74.11 92.63
2HSL [galdran2023multi] 82.90 - -
OUSM [xue2019robust] 80.53 73.70 93.69
SELF [nguyen2019self] 80.49 75.31 93.99
DiRL [kapse2024attention] 83.60 - 95.01
SH-PEFT [liu2024sparsity] 84.80 80.61 -
UNI [chen2024towards] 85.51 85.29 95.21
PolarHE (ours) 86.70 86.42 95.96
Patch Classification by Linear Probing. To evaluate the quality of the representations learned by PolarHE, we adopt a linear probing protocol, where a lightweight classifier is trained on top of the frozen encoder. This setup, following standard benchmarks [kang2023benchmarking], isolates the representation quality from task-specific fine-tuning.
We report the average performance over three random seeds on both the Chaoyang and MHIST datasets to ensure result robustness. For fair benchmarking, we compare against results reported in prior work or reproduced using publicly available code and official configurations.
Our method, PolarHE, achieves state-of-the-art performance on both datasets. On the Chaoyang Dataset from Table I, it achieves an accuracy of 86.70%, an F1 score of 86.42%, and an AUC of 95.96%. On the MHIST Dataset from Table II, it achieves an accuracy of 89.06%, an F1 score of 88.97%, and an AUC of 94.10%.
TABLE II: Test performance on MHIST [wei2021petri] dataset.
Method Acc F1 AUC
DINO [caron2021emerging] 83.10 81.60 89.78
BarlowTwins [zbontar2021barlow] 81.27 - -
MoCo-v2 [kang2023benchmarking] 85.88 - -
Bitfit [zaken2021bitfit] 82.40 80.86 89.60
SSF [lian2022scaling] 82.29 80.85 89.29
RMSGD [hosseini2022exploiting] 82.58 - -
TPP [lei2024pre] 83.42 82.21 91.38
DiRL [kapse2024attention] 78.20 - 87.11
Virchow [vorontsov2023virchow] 83.40 83.51 -
CTransPath [wang2021transpath] 83.75 - -
UNI [chen2024towards] 85.63 85.57 91.98
PolarHE (ours) 89.06 88.97 94.10
Importantly, only H&E images are used during both linear probing and inference. This demonstrates that our polarization-guided pretraining substantially enhances the H&E encoder’s downstream performance—even when polarization data is unavailable at test time. This reflects real-world deployment settings, where polarization imaging is costly or limited in availability, yet its benefits can still be leveraged through training-time supervision.
@@
Ablation Study. Table III shows removing intra-modal training or modality decoupling reduces accuracy, confirming their importance. Without intra-modal training, performance drops, especially on Chaoyang, highlighting its role in reinforcing modality-specific features. Without modality decoupling, accuracy also declines, showing the need for feature separation. The full model performs best, validating both components. For common representation ratio in Table IV, the downstream performance increases and decreases smoothly along with the change of the percentage of common dimensions, which indicates the sparsity of the common embedding space.
TABLE III: Ablation study on loss components. We report the accuracy on Chaoyang and MHIST datasets.
@@
TABLE IV: Effect of common representation ratio on accuracy.
@@
V Conclusion
We propose PolarHE, a dual-modality fusion framework that integrates polarization imaging with H&E staining for enhanced representation learning in pathology. By leveraging a feature decomposition strategy, our approach effectively disentangles common and modality-specific features, leading to improved multimodal representation learning. Through comprehensive experiments, we demonstrated that the integration of H&E and polarization representations leads to significantly better performance compared to existing methods. These findings underscore the untapped potential of polarization imaging as a powerful auxiliary modality in computational pathology. By enhancing diagnostic accuracy and model robustness, PolarHE marks a promising step toward more interpretable and automated pathology analysis, paving the way for advancements in multimodal deep learning for clinical applications.
VI Acknowledgment
[…]